OpenStudy (anonymous):

Find the zeros of: y = 2x^3 + 2x^2 - 40x.

OpenStudy (bgrg007):

set the value of y to 0 and then solve. 0 = 2x^3 + 2x^2 - 40x factor out the x 0 = x(2x^2 + 2x - 40) then the 2 to make it easier 0 = 2x(x^2+x-20) 0 = 2x(x-4)(x+5) so x = 0, -5, 4 <- answer

OpenStudy (anonymous):

@BGrg007 - thank you so much!

OpenStudy (bgrg007):

np